Find the positive roots of the equation 4x squared +27x=40 by completing the square
Answer by stanbon(75887)

Find the positive roots of the equation 4x squared +27x=40 by completing the square
------------
4x^2 + 27x = 40
---
4(x^2+(27/4)x + (27/8)^2) = 40+ 4(27/8)^2
-----
4(x+(27/8))^2 = 85.5625
---
(x+(27/8)) = +/-sqrt[85.5625/4]
---
x+(27/8) = +/-4.625
---
x = -(27/8)+-4.625
========================
x = 1.25 and x = -8
============================
Cheers,
Stan H.